OpenStudy (anonymous):

etermine which statement is false. |1 + 5| = |1| + |5| |8 − 3| = |8| − |3| |7 + 3| = |7| + |3| |3 − 6| = |3| − |6|

OpenStudy (chriss):

the fourth one

OpenStudy (unklerhaukus):

|3 − 6| = |3| − |6| |-3| = 3-6 3 = -3 which is false